# Black hole spectroscopy of collapsing and merging neutron stars

### v1

@@ IV.1 Ringdown models
The alternative approach for modeling the same ringdown signal Eq. (5) is called ”theory-agnostic”. Within this method, each $\omega_{\ell mn}$ will be described by an independent real and imaginary part. In the theory-agnostic case, the total number of free parameters is $4\times N$ (where N is the number of included QNMs). In the theory-specific case, it is only $2\times N+2$, which has profound impact on the stability of the fitting procedure when investigating a larger number of QNMs.
@@ IV.3 Fitting routine and the mismatch
where $y_{i}$ is the $i$-th data and $t_{i}$ the $i$-th time point, and $h$ is our ringdown model in Eq. (5) together with the fitted parameters. curve_fit can take the error on the data as an argument. For the systems considered, we constrained the error in the data to about $10^{-7}$, which was usually the onset of the numerical noise floor. Although this error is not related to a measurement, it provides a qualitative estimate of how numerical errors might propagate to the fitted parameters. On the other hand, the error on the fitted parameters can be computed via
@@ V Applications
In the following, we analyze the ringdown signal of the Class I model C8 in Sec. V.1. We then provide a more condensed summary of the ringdown analysis for the other Class I models in Sec. V.2. The early part of the Class II model B5 is investigated in Sec. V.3. The ringdown analysis of our three selected BNS simulations are shown in Sec. V.4-V.6.
@@ V.1 QNMs of simulation C8
Figure 10: The difference of the C8 theory-specific fits for our 3 ringdown models (consult Sec. IV) once with a freely varying spin and the other time for a fixed spin. Upper panel: mismatches for all cases, as indicated in the legend. Lower panel: difference between the BH mass from the simulation and the fitted mass, defined in Eq. (6). Errors are only drawn as bands around the fits with fixed spin.
@@ V.2 QNMs of combined Class I simulations
In the following, we provide a condensed analysis of all 53 Class I models. We limit the length of the ringdown signal to $110$ after the peak, so that we fit the frequencies in the same time window for all cases, and get a more coherent picture. The fits are performed for each model at 60 different starting times in the range from 0 to $60$, to leave enough signal for the later starting times. To explore the possibility that non-vacuum effects impact the fundamental mode, we report the corresponding baryonic mass outside the BH as a qualitative measure.
@@
Figure 11: Percentage deviations of $\delta\omega_{R}$ in log scale against the baryonic mass outside the BH $\delta m$ scaled by the BH mass $M_{\mathrm{BH}}$. The legend indicates markers for the QNM fit at different starting times after the peak. Each marker is connected by a grey dotted line for each model to trace the evolution of the frequency deviation. A thin black line is drawn to indicate a threshold of 1%. Top: ringdown model with just the fundamental mode $(N=1)$. Middle: ringdown model with the fundamental mode and first overtone $(N=2)$. Bottom: ringdown model with fundamental mode and the first two overtones $(N=3)$.
@@ V.3 Early B5 QNM spectrum
The combined theory-agnostic and -specific fits are shown in Fig. 14. As expected, the mismatches decrease when we increase the number of modes in the ringdown model. For $N=2$ and $N=3$, the mismatch approaches that of $N=1$ near the end of the initial time range, indicating that the modeled overtones reach their fitting bounds and do not improve the signal anymore (more details can be found in Appendix B; we find a stable fundamental mode, while only observing a semi-stable first overtone for the first part of the analysis). We observe that for $N=1$, the mismatch between the vacuum and the fitted values for $\mathrm{Re}(\omega)$ decreases significantly once we move to later starting times, approaching around 1% when fitting 150 time steps after the peak of the signal. Including additional modes leads to a 1% difference even at early times. The deviation of $\mathrm{Im}(\omega)$ from the predictions closely resembles that of the real part, and is surprisingly accurate at late times. Mass and spin extractions show us a similar picture, with small deviations compared to the true values. Especially at early times, the fitting procedure seems to have trouble converging to the true spin, which improves at later times (this can be best seen in Fig. 19 of Appendix B, where we show the starting time evolution in the MaM-a plane).
@@
Lastly, the ringdown analysis of the lightest of the three BNS simulations, i.e., the $M_{\mathrm{tot}}=2.7M_{\odot}$ binary, is shown in Fig. 16. Due to the shortness of the signal, we cannot extend the length of the signal to longer than $150$ after BH formation. To consider enough data points at later starting times, we set the maximum starting time to $60$ after AH formation. Starting with the theory-agnostic fits, we see that for the $N=1$ ringdown model, deviations to the prediction of $\mathrm{Re}(\omega)$ are around 5% for most of the signal. This, of course, decreases once we consider more modes. For the $N=2$ and $N=3$ models, this reduces to about 1% difference from the vacuum values, either due to the baryonic matter left outside or modeling specifications. The imaginary part for all ringdown models is above the 1% threshold for most starting times. In general the fits are not as ”clean” as for the $M_{\mathrm{tot}}=3.8M_{\odot}$ system, since most likely the baryonic matter outside affects the ringdown morphology which influences the parameter extraction. This is also visible, when looking at the theory-specific results in the bottom panel of Fig. 16. At early times for the multi-mode models and late times for the one-mode model, we see deviations of 3-4% for both mass and spin. With $\sim 6.5$ kHz the fundamental ringdown frequency is in the high-frequency regime making it unlikely to be detected.
@@ VI Conclusions
In traditional BH spectroscopy, one utilizes Teukolsky’s equation [28] to investigate gravitational perturbations of a Kerr spacetime in vacuum [22]. The ringdown of BBH mergers is described by this well-established assumption, but BNS and BHNS mergers violate it. Assessing the impact of dynamical matter environments on the ringdown is, in general, non-trivial. Current GW detectors are not able to measure QNMs from collapsing NSs in such systems, mainly because their QNM frequencies are outside the sensitivity band. However, next-generation GW detectors like the Einstein Telescope [120, 121, 119] and Cosmic Explorer [118, 122] might enable precision measurements of the ringdown of such systems, although the ringdown will most likely be seen in the less sensitive high-frequency regime.
@@
Signals of Class I are resembling a typical BH ringdown, but could in principle be prone to non-vacuum effects, which might be interpreted as false violations from GR. We therefore focused on using standard methods in BH spectroscopy to analyze such signals from collapsing isolated NSs and for BNS mergers in Sec. V. Whether matter matters is highly case-dependent. Even though they are unstable and physically questionable, adding more QNMs to our ringdown models significantly affects the accuracy of parameter estimation for important properties such as the remnant mass and its fundamental mode. Matter clearly matters for systems of Class II and Class III as they show clear deviations from a typical ringdown signal. Class II ringdowns admit qualitative similarities to cases presented in Ref. [60], which studied oscillations of a matter shell surrounding a BH. Nevertheless, it is improbable that any confusion with a BH ringdown will occur, and if they are observed, they would not be employed for conventional BH spectroscopy. Phenomenological ringdown models may be used to extract disk properties for these systems [61].
In our study, we find that using only the dominant QNM for the ringdown is not sufficient to obtain the correct BH mass below a percent level threshold. It appears that the presence of matter are not the causes of these deviations, at least not for waveforms of Class I. By adding one or two overtones, mismatches are greatly reduced, and the reconstructed BH mass is often even accurate to within a few percent. Furthermore, it provides more consistent values for the amplitude and phase of the fundamental mode, albeit not as consistent as in the vacuum case or linear simulations.
@@
There are some aspects of our work that could be extended in the future. Our analysis uses a least-squares-minimization routine approximating the errors of the data points (that are always (implicitly) chosen in such an analysis), with those of the numerical-relativity noise floor. GW data analysis pipelines could be utilized to incorporate our numerical-relativity simulations as an injection. Such an analysis would yield measurement uncertainties regarding the ringdown parameters, which would be contingent upon the detector specification. Moreover, this would make it possible to measure the ratio of statistical to systematic errors, a crucial metric for determining the suitability of the underlying model used to analyze a signal. Using linear-signal-analysis (LSA) [123, 124, 125, 126], in particular the bias formula [125, 126], the bias ratio could also be explored by using the numerical-relativity simulations as an injection, and including more QNMs in a ringdown model, until systematic effects are smaller than statistical ones. This has recently been explored in Ref. [127] in the vacuum case, while the range of validity of LSA predictions for ringdown systematics has been explored in Ref. [128] by explicitly comparing it to Bayesian analysis using Mark-chain-Monte-Carlo sampling.
@@ Acknowledgements.
O. S. thanks Anna Neuweiler for help regarding the creation of BNS initial data with SGRID and their evolution with BAM, as well as Ivan Markin for useful discussions related to issues with BAM. Additionally, O. S. thanks Henrique Gieg and Anna Puecher for help with computations on signal-to-noise ratio estimates. S. H. V. thanks Arnab Dhani for useful discussions related to the detectability of our systems with next-generation GW detectors. S. H. V. acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG): Project No. 386119226. T. D. acknowledges funding from the European Union (ERC, SMArt, 101076369). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Computations for this paper have been performed on the DFG-funded research cluster Jarvis at the University of Potsdam (INST 336/173-1; project number: 502227537).
@@
Data on the waveforms, baryonic masses, and black hole properties from the horizon finder for the simulations discussed in this work have been made available under https://doi.org/10.5281/zenodo.16792099. Additional data are available from the authors upon reasonable request.
@@ Appendix A Isolated Neutron Star Equilibrium Models
In the following, we list the parameters of all 130 dRNS models, which were created with the procedure in Sec. II.1.1. The parameters are dimensionless, based on the rescalings of [78, 79]. Each model in the table obtains an identifier according to the following procedure: 1. The first half of the identifier is a letter corresponding to the respective rotation parameter of the model, e.g., the letter $A$ for $\hat{A}=1.0$, the letter $B$ for $\hat{A}=0.9$, all the way to $J$ for $\hat{A}=0.1$. 2. Next to the letter, we give the model a number based on its rotation parameter, e.g., the number 1 for $r_{p}/r_{e}=0.35$, the number 2 for $r_{p}/r_{e}=0.4$, up to the number 13 for $r_{p}/r_{e}=0.95$. We group all models with the same rotation parameter under a Series with the letter used in the model name, i.e., the models with $\hat{A}=1.0$ under Series A, the models with $\hat{A}=0.9$ under Series B, and so on. Next to the model identifier, we show the respective class based on the classification scheme presented in Sec. III. The classes are represented in different colors to better distinguish them: green for Class I, orange for Class II, and red for Class III. Some models are denoted with an asterisk, which indicates that for this particular model the central density had to be increased to induce the collapse of the NS. The adjusted value is given in the $\rho_{c}$ column behind the interpolated value from the corresponding model sequence. A double asterisk is given behind two model identifiers because these models showed problematic output for the baryonic mass around the remnant. It seemed as if the baryonic mass increased to its pre-collapse value and then went down again. To prevent the contamination of these results to our overall analysis, we re-ran these two models at a different central density to prevent the issue from occurring again.
@@ Appendix B Mode stability and starting time dependence
Figure 18: Similar to 17 for the $M_{\mathrm{tot}}=3.8M_{\odot}$ BNS system. In the middle column, for the first overtone, we show a grey shaded region to indicate the rough onset of instability of the mode.
@@
The mode stability depends generally on the studied system. Here, we start with a stability analysis of the Class I model C8, shown in Fig. 17. Since we consider three different ringdown models with $N=1,2,3$ modes, we can examine the stability of the fundamental mode and the first two overtones if used in the respective model. We observe a stable fundamental mode amplitude and phase across all ringdown models, where stability improves upon the inclusion of additional modes. The same cannot be said about the first and second overtone which are quite unstable for the multi-mode models. This suggests something already pointed out in the literature for the analysis of vacuum simulations [43, 44, 45, 46, 50], namely, that adding more modes can drastically reduce the mismatches between parameters, even if these modes might be physically unrealistic or even absent. A similar behavior is observed across all Class I ringdowns.
@@
Fig. 18 shows the mode stability for the $M_{\mathrm{tot}}=3.8M_{\odot}$ BNS. Again, the fundamental mode amplitude and phase are stable, although the phase is quite off at the beginning of the $N=1$ ringdown model, which improves for the multi-mode models. However, at least for the first overtone, we can observe a semi-stable amplitude and phase for the first half of the starting time window. Afterward, the amplitude and phase grow in value and at some point hit their respective bounds. So at least for the first half of the ringdown, we can claim that a first overtone might be physically present, which is not the case for the other two BNS cases. But when looking at $\delta\omega_{R}$ and $\delta\omega_{I}$, we observe deviations of $\sim 10\%$ and $\gtrsim 30\%$, respectively, which weakens the argument. On the other hand, the second overtone seems to be unstable over the entire range with wrong values for the frequency and damping time. Also, the phase of the fundamental mode and the first overtone should be phase anti-aligned by a factor of $\pi$, which is something we could at least observe for the $3.8M_{\odot}$ binary; see Fig. 18. In Ref. [128], it was suggested, that the bias of inferred parameters should oscillate with the fundamental mode frequency. We found that this is indeed the case for some unmodeled effects when one considers the amplitude subtracted by the mean of its late-time value.
@@
In addition to our discussion on mode stability, we investigated the starting time evolution for the theory-agnostic and -specific fits of the three selected BNS ringdown signals. The top panels of Fig. 19 show the theory-agnostic results in the $\mathrm{Re}(\omega)-\mathrm{Im}(\omega)$ plane for the three ringdown models used. When considering only the fundamental mode, we can see a cycloidal pattern across the starting time range for the heavy and light BNS. While the heavy $M_{\mathrm{tot}}=3.8M_{\odot}$ binary converges to the Kerr predictions at late times, the other systems are further from the vacuum frequencies, even at late starting times. Convergence for the heavy BNS happens faster when considering two or three modes. While this is also the case for the other two systems, they still do not converge to the predictions, but there is a visible offset. A similar picture emerges, when looking at the starting time evolution of the theory-specific fits in the MaM-a plane, shown in the bottom panels of Fig. 19. The cycloidal pattern is also visible for some systems when using a multi-mode ringdown model. As before, including additional modes improves the convergence, although the $M_{\mathrm{tot}}=2.7M_{\odot}$ and $M_{\mathrm{tot}}=3.2M_{\odot}$ systems do not convergence to the “true” mass and spin as determined by the simulation.

### v2

@@ IV.1 Ringdown models
The alternative approach for modeling the same ringdown signal Eq. (5) is called “theory-agnostic”. Within this method, each $\omega_{\ell mn}$ will be described by an independent real and imaginary part. In the theory-agnostic case, the total number of free parameters is $4\times N$ (where N is the number of included QNMs). In the theory-specific case, it is only $2\times N+2$, which has profound impact on the stability of the fitting procedure when investigating a larger number of QNMs.
@@ IV.3 Fitting routine and the mismatch
where $y_{i}$ is the $i$th data and $t_{i}$ the $i$th time point, and $h$ is our ringdown model in Eq. (5) together with the fitted parameters. curve_fit can take the error on the data as an argument. For the systems considered, we constrained the error in the data to about $10^{-7}$, which was usually the onset of the numerical noise floor. Although this error is not related to a measurement, it provides a qualitative estimate of how numerical errors might propagate to the fitted parameters. On the other hand, the error on the fitted parameters can be computed via
@@ V Applications
In the following, we analyze the ringdown signal of the Class I model C8 in Sec. V.1. We then provide a more condensed summary of the ringdown analysis for the other Class I models in Sec. V.2. The early part of the Class II model B5 is investigated in Sec. V.3. The ringdown analysis of our three selected BNS simulations are shown in Secs. V.4-V.6.
@@ V.1 QNMs of simulation C8
Figure 10: The difference of the C8 theory-specific fits for our three ringdown models (consult Sec. IV) once with a freely varying spin and the other time for a fixed spin. Upper panel: Mismatches for all cases, as indicated in the legend. Lower panel: Difference between the BH mass from the simulation and the fitted mass, defined in Eq. (6). Errors are only drawn as bands around the fits with fixed spin.
@@ V.2 QNMs of combined Class I simulations
In the following, we provide a condensed analysis of all 53 Class I models. We limit the length of the ringdown signal to $110$ after the peak, so that we fit the frequencies in the same time window for all cases, and get a more coherent picture. The fits are performed for each model at 60 different starting times in the range from 0 to $60$, to leave enough signal for the later starting times. To explore the possibility that nonvacuum effects impact the fundamental mode, we report the corresponding baryonic mass outside the BH as a qualitative measure.
@@
Figure 11: Percentage deviations of $\delta\omega_{R}$ in log scale against the baryonic mass outside the BH $\delta m$ scaled by the BH mass $M_{\mathrm{BH}}$. The legend indicates markers for the QNM fit at different starting times after the peak. Each marker is connected by a gray dotted line for each model to trace the evolution of the frequency deviation. A thin black line is drawn to indicate a threshold of 1%. Top: Ringdown model with just the fundamental mode $(N=1)$. Middle: Ringdown model with the fundamental mode and first overtone $(N=2)$. Bottom: Ringdown model with fundamental mode and the first two overtones $(N=3)$.
@@ V.3 Early B5 QNM spectrum
The combined theory-agnostic and -specific fits are shown in Fig. 14. As expected, the mismatches decrease when we increase the number of modes in the ringdown model. For $N=2$ and $N=3$, the mismatch approaches that of $N=1$ near the end of the initial time range, indicating that the modeled overtones reach their fitting bounds and do not improve the signal anymore (more details can be found in Appendix B; we find a stable fundamental mode, while only observing a semistable first overtone for the first part of the analysis). We observe that for $N=1$, the mismatch between the vacuum and the fitted values for $\mathrm{Re}(\omega)$ decreases significantly once we move to later starting times, approaching around 1% when fitting 150 time steps after the peak of the signal. Including additional modes leads to a 1% difference even at early times. The deviation of $\mathrm{Im}(\omega)$ from the predictions closely resembles that of the real part, and is surprisingly accurate at late times. Mass and spin extractions show us a similar picture, with small deviations compared to the true values. Especially at early times, the fitting procedure seems to have trouble converging to the true spin, which improves at later times (this can be best seen in Fig. 19 of Appendix B, where we show the starting time evolution in the MaM-a plane).
@@
Lastly, the ringdown analysis of the lightest of the three BNS simulations, i.e., the $M_{\mathrm{tot}}=2.7M_{\odot}$ binary, is shown in Fig. 16. Due to the shortness of the signal, we cannot extend the length of the signal to longer than $150$ after BH formation. To consider enough data points at later starting times, we set the maximum starting time to $60$ after AH formation. Starting with the theory-agnostic fits, we see that for the $N=1$ ringdown model, deviations to the prediction of $\mathrm{Re}(\omega)$ are around 5% for most of the signal. This, of course, decreases once we consider more modes. For the $N=2$ and $N=3$ models, this reduces to about 1% difference from the vacuum values, either due to the baryonic matter left outside or modeling specifications. The imaginary part for all ringdown models is above the 1% threshold for most starting times. In general the fits are not as ”clean” as for the $M_{\mathrm{tot}}=3.8M_{\odot}$ system, since most likely the baryonic matter outside affects the ringdown morphology which influences the parameter extraction. This is also visible, when looking at the theory-specific results in the bottom panel of Fig. 16. At early times for the multimode models and late times for the one-mode model, we see deviations of 3-4% for both mass and spin. With $\sim 6.5$ kHz the fundamental ringdown frequency is in the high-frequency regime making it unlikely to be detected.
@@ VI Conclusions
In traditional BH spectroscopy, one utilizes Teukolsky’s equation [28] to investigate gravitational perturbations of a Kerr spacetime in vacuum [22]. The ringdown of BBH mergers is described by this well-established assumption, but BNS and BHNS mergers violate it. Assessing the impact of dynamical matter environments on the ringdown is, in general, nontrivial. Current GW detectors are not able to measure QNMs from collapsing NSs in such systems, mainly because their QNM frequencies are outside the sensitivity band. However, next-generation GW detectors like the Einstein Telescope [124, 125, 123] and Cosmic Explorer [122, 126] might enable precision measurements of the ringdown of such systems, although the ringdown will most likely be seen in the less sensitive high-frequency regime.
@@
Signals of Class I are resembling a typical BH ringdown, but could in principle be prone to nonvacuum effects, which might be interpreted as false violations from GR. We therefore focused on using standard methods in BH spectroscopy to analyze such signals from collapsing isolated NSs and for BNS mergers in Sec. V. Whether matter matters is highly case dependent. Even though they are unstable and physically questionable, adding more QNMs to our ringdown models significantly affects the accuracy of parameter estimation for important properties such as the remnant mass and its fundamental mode. Matter clearly matters for systems of Class II and Class III as they show clear deviations from a typical ringdown signal. Class II ringdowns admit qualitative similarities to cases presented in Ref. [60], which studied oscillations of a matter shell surrounding a BH. Nevertheless, it is improbable that any confusion with a BH ringdown will occur, and if they are observed, they would not be employed for conventional BH spectroscopy. Phenomenological ringdown models may be used to extract disk properties for these systems [61].
In our study, we find that using only the dominant QNM for the ringdown is not sufficient to obtain the correct BH mass below a percent level threshold. It appears that the presence of matter are not the causes of these deviations, at least not for waveforms of Class I. By adding one or two overtones, mismatches are greatly reduced, and the reconstructed BH mass is often even accurate to within a few percent. This underlines that adding overtones does not only provide better fits, but moreover, improves the extraction of physical parameters. Furthermore, it provides more consistent values for the amplitude and phase of the fundamental mode, albeit not as consistent as in the vacuum case or linear simulations.
@@
There are some aspects of our work that could be extended in the future. Our analysis uses a least-squares minimization routine approximating the errors of the data points (that are always (implicitly) chosen in such an analysis), with those of the numerical-relativity noise floor. Bayesian analysis has been applied to real data in the context of determining overtones, which shows how the unclear starting time and impact of noise influence the analysis, see Refs. [127, 128, 129, 130] as an example discussing the difficulties. GW data analysis pipelines could be utilized to incorporate our numerical-relativity simulations as an injection, e.g., similar to Ref. [131], which investigates the detectability of black hole formation through BNS mergers with next generation detectors. Such an analysis would also yield measurement uncertainties regarding the ringdown parameters, which would be contingent upon the detector specification. Moreover, this would make it possible to measure the ratio of statistical to systematic errors, a crucial metric for determining the suitability of the underlying model used to analyze a signal. Using linear-signal analysis (LSA) [132, 133, 134, 135], in particular the bias formula [134, 135], the bias ratio could also be explored by using the numerical-relativity simulations as an injection, and including more QNMs in a ringdown model, until systematic effects are smaller than statistical ones. This has recently been explored in Ref. [136] in the vacuum case, while the range of validity of LSA predictions for ringdown systematics has been explored in Ref. [137] by explicitly comparing it to Bayesian analysis using Mark-chain-Monte-Carlo sampling.
@@ Acknowledgements.
O. S. thanks Anna Neuweiler for help regarding the creation of BNS initial data with SGRID and their evolution with BAM, as well as Ivan Markin for useful discussions related to issues with BAM. Additionally, O. S. thanks Henrique Gieg and Anna Puecher for help with computations on signal-to-noise ratio estimates. S. H. V. thanks Arnab Dhani for useful discussions related to the detectability of our systems with next-generation GW detectors. S. H. V. acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG): Project No. 386119226. T. D. acknowledges funding from the European Union (ERC, SMArt, 101076369). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Computations for this paper have been performed on the DFG-funded research cluster Jarvis at the University of Potsdam (INST 336/173-1; Proj. No.: 502227537).
@@
The data that support the findings of this article are openly available [138].
@@
In the following, we list the parameters of all 130 dRNS models, which were created with the procedure in Sec. II.1.1. The parameters are dimensionless, based on the rescalings of [79, 80]. Each model in the table obtains an identifier according to the following procedure: 1. The first half of the identifier is a letter corresponding to the respective rotation parameter of the model, e.g., the letter $A$ for $\hat{A}=1.0$, the letter $B$ for $\hat{A}=0.9$, all the way to $J$ for $\hat{A}=0.1$. 2. Next to the letter, we give the model a number based on its rotation parameter, e.g., the number 1 for $r_{p}/r_{e}=0.35$, the number 2 for $r_{p}/r_{e}=0.4$, up to the number 13 for $r_{p}/r_{e}=0.95$. We group all models with the same rotation parameter under a Series with the letter used in the model name, i.e., the models with $\hat{A}=1.0$ under Series A, the models with $\hat{A}=0.9$ under Series B, and so on. Next to the model identifier, we show the respective class based on the classification scheme presented in Sec. III. Some models are denoted with an asterisk, which indicates that for this particular model the central density had to be increased to induce the collapse of the NS. The adjusted value is given in the $\rho_{c}$ column behind the interpolated value from the corresponding model sequence. A double asterisk is given behind two model identifiers because these models showed problematic output for the baryonic mass around the remnant. It seemed as if the baryonic mass increased to its precollapse value and then went down again. To prevent the contamination of these results to our overall analysis, we reran these two models at a different central density to prevent the issue from occurring again.
@@
Figure 18: Similar to 17 for the $M_{\mathrm{tot}}=3.8M_{\odot}$ BNS system. In the middle column, for the first overtone, we show a gray shaded region to indicate the rough onset of instability of the mode.
@@
The mode stability depends generally on the studied system. Here, we start with a stability analysis of the Class I model C8, shown in Fig. 17. Since we consider three different ringdown models with $N=1,2,3$ modes, we can examine the stability of the fundamental mode and the first two overtones if used in the respective model. We observe a stable fundamental mode amplitude and phase across all ringdown models, where stability improves upon the inclusion of additional modes. The same cannot be said about the first and second overtone which are quite unstable for the multimode models. In particular, from Fig. 17 it is apparent, that partial stability is only present at the start of the ringdown signal. Afterwards, the amplitude and phase of the first and second overtone blow up and hit the respective boundaries set by the prior. Increasing the bounds does not help with stability, but only delays the point when the bounds are hit. The real and imaginary part of the mode frequency on the other hand vary largely and can easily be off by more than 30%. Other fitting algorithms might be able to extract overtones that are more stable, although these could be artificial. As there is no applicable ringdown model that treats the nonlinear effects better, overtones remain hard to extract. This suggests something already pointed out in the literature for the analysis of vacuum simulations [43, 44, 45, 46, 50], namely, that adding more modes can drastically reduce the mismatches between parameters, even if these modes might be physically unrealistic or even absent. A similar behavior is observed across all Class I ringdowns.
@@
Figure 18 shows the mode stability for the $M_{\mathrm{tot}}=3.8M_{\odot}$ BNS. Again, the fundamental mode amplitude and phase are stable, although the phase is quite off at the beginning of the $N=1$ ringdown model, which improves for the multimode models. However, at least for the first overtone, we can observe a semistable amplitude and phase for the first half of the starting time window. Afterward, the amplitude and phase grow in value and at some point hit their respective bounds. So at least for the first half of the ringdown, we can claim that a first overtone might be physically present, which is not the case for the other two BNS cases. But when looking at $\delta\omega_{R}$ and $\delta\omega_{I}$, we observe deviations of $\sim 10\%$ and $\gtrsim 30\%$, respectively, which weakens the argument. On the other hand, the second overtone seems to be unstable over the entire range with wrong values for the frequency and damping time. These results are in agreement with the above discussion for our Class I model. Also, the phase of the fundamental mode and the first overtone should be phase anti-aligned by a factor of $\pi$, which is something we could at least observe for the $3.8M_{\odot}$ binary; see Fig. 18. In Ref. [137], it was suggested, that the bias of inferred parameters should oscillate with the fundamental mode frequency. We found that this is indeed the case for some unmodeled effects when one considers the amplitude subtracted by the mean of its late-time value.
@@
In addition to our discussion on mode stability, we investigated the starting time evolution for the theory-agnostic and -specific fits of the three selected BNS ringdown signals. The top panels of Fig. 19 show the theory-agnostic results in the $\mathrm{Re}(\omega)-\mathrm{Im}(\omega)$ plane for the three ringdown models used. When considering only the fundamental mode, we can see a cycloidal pattern across the starting time range for the heavy and light BNS. While the heavy $M_{\mathrm{tot}}=3.8M_{\odot}$ binary converges to the Kerr predictions at late times, the other systems are further from the vacuum frequencies, even at late starting times. Convergence for the heavy BNS happens faster when considering two or three modes. While this is also the case for the other two systems, they still do not converge to the predictions, but there is a visible offset. A similar picture emerges, when looking at the starting time evolution of the theory-specific fits in the MaM-a plane, shown in the bottom panels of Fig. 19. The cycloidal pattern is also visible for some systems when using a multimode ringdown model. As before, including additional modes improves the convergence, although the $M_{\mathrm{tot}}=2.7M_{\odot}$ and $M_{\mathrm{tot}}=3.2M_{\odot}$ systems do not convergence to the “true” mass and spin as determined by the simulation.